### 30 Jan 2024 v1

@@ 2.2 IRAM 30-meter observations of Orion B
We use spectroscopic data of cyanide (CN) and formyl cation (HCO) ( transition line) to trace the UV-irradiated gas (Bron et al., 2018). CN is present in UV-illuminated edges as a photodissociation product of HCN and HNC, resulting from UV-dominated chemistry, and can be collisionally excited by electrons and neutrals (Santa-Maria et al., 2023). Therefore, the CN rotational lines remain bright in dense UV-illuminated edges, which makes them excellent tracers of the UV-dominated regions, like the one seen in NGC 2024. In addition, CN shows the multiple hyperfine components (Penzias et al., 1974), spectrally resolved by our observations. We report the properties of each hyperfine component used in this work in Tab. 1, which allow derivation of the opacity (Milam et al., 2009) and excitation temperature of the (1–0) transition, enabling accurately constraining its excitation and obtaining limits on the gas density.
HCO , similarly to CN, traces photon-dominated regions (Young Owl et al., 2000a; Lis & Schilke, 2003). HCO has different production pathways dominant at different physical conditions of the gas. At the edges of the dissociation region, HCO is a precursor of CO, and can be produced through the CO (Hogerheijde et al., 1995; Young Owl et al., 2000b; van der Werf et al., 1996) or CH molecule (e.g., Goicoechea et al., 2016). Recent studies detected HCO emission the edge of the PDR in the Orion bar (Goicoechea et al., 2016), and the Horsehead nebula (Hernández-Vera et al., 2023). Pety et al. (2017); Gratier et al. (2017) found that CN and HCO are sensitive to the UV radiation. Moreover, by applying a clustering algorithm, Bron et al. (2018) found that CN and HCO trace the UV-radiated gas, contrary to the CO molecule, which gets photodissociated. Therefore, using CN and HCO emission in our work, together with the FUV heated dust traced by SOFIA HAWC+ observations, we can directly assess the gas impacted by the radiative stellar feedback in NGC 2024.
@@ 3 Dust polarization
The dust continuum emission is strongest towards the center of NGC 2024. A significant emission is observed along the filament that is in the front of the ionization region caused by the young massive star IRS 2b (the central part of our map, Bik et al., 2003), and along the ionization fronts on the southeastern and southwestern part of NGC 2024 (labeled as Bubble West and Bubble East, and shown in the middle row in Fig. 2).
@@
In addition, we compare our magnetic field directions to the direction inferred from the near-infrared (NIR) polarization of young stars (Kandori et al., 2007). The NIR polarization directly traces the direction of the magnetic field (i.e., there is no difference as for the FIR dust polarization - Sec. 2.1). We show this comparison in the top panels of Fig. 11 in App. A.2. As seen from the figure, we find that magnetic field vectors derived from NIR polarization and our work are parallel to each other, indicating an agreement between these two datasets in NGC 2024 (see Fig. 11 in App. A.2). In addition, we compare the magnetic field direction with the 100 µm (Dotson et al., 2000) and sub-millimeter (850 µm) polarization measurements (Matthews et al., 2002) in the central part of NGC 2024 (labeled with the dashed black rectangle in top panels of Fig. 11) and show zoomed-in panels in the bottom row of Fig. 11.
@@
The area labeled with dark blue contours in the top panels in Fig. 11 is particularly interesting because we observe an outflowing feature in the HCO emission at velocities of km s. The outflow originates from the FIR 5 source located in the dense molecular cloud behind the expanding HII region (e.g., Richer et al., 1992; Greaves et al., 2001; Choi et al., 2015). Dust continuum emission at 154 µm possibly traces the outflow in this region (see also the bottom middle panel in Fig. 3), as we note that magnetic field lines follow the direction of the outflow. We have not observed this behavior at 214 µm.
@@
The magnetic flux is frozen within the molecular gas, and as a consequence, the -field will trace its morphology. Therefore, the local environmental conditions that shape the distribution of molecular gas will also impact the magnetic field. On the one hand, the magnetic field is highly ordered in some regions of NGC 2024; for instance, -field follows the dusty filament to the south of NGC 2024 (see, e.g., the bottom left panel of Fig. 2), except for the very southern part of the filament (close to the bottom left corner of the top panel in Fig. 2), where the magnetic field is perpendicular to the filament. At the northern part of NGC 2024, the magnetic field direction varies from following the filament to being perpendicular.
@@
We observe the nearly horizontal direction of the -field at the edges of the expanding HII region (middle panels in Fig. 2). Gas affected by the stellar feedback (e.g., UV radiation, stellar winds) is pushed outward from the HII region. Consequently, magnetic field lines become parallel to the edges of such an expanding shell (e.g., Tahani et al., 2023). On the other hand, the magnetic field appears chaotic in the central and densest parts of NGC 2024. This area shows great complexity, as it results from the mixture of the filament located in the front, the HII region and the dense molecular gas in the background (see, e.g., Fig. 8 in Matthews et al. 2002, or Fig. 5 in Roshi et al. 2014). One possible explanation for this observed morphology of the magnetic field is that the magnetic field changes its orientation relative to the filament depending on the column density contrast between the filament and the background emission, from being parallel to perpendicular to the molecular gas distribution ( cm, e.g., Planck Collaboration et al., 2016c; Alina et al., 2019). Additionally, the magnetic field could be ”pinched” due to the gravitational collapse of the gas (Basu, 2000; Lai et al., 2002; Doi et al., 2020, 2021), which can also explain the magnetic field direction at the northern part of NGC 2024 and the very southern part of the filament.
@@ 4.1 Moment maps
Fig. 3 shows moment maps for the CN(1–0) (top row) and HCO(1–0) (bottom row) emission. We create these moment maps using CUBE in GILDAS. Before creating moment maps, we mask CN and HCO cubes. The mask is created using the segmentation technique in CUBE, where we identify neighboring voxels with continuous S/N ratio; we have selected all voxels with S/N ratio above 2 (see Einig et al., 2023, for a complete description). In this case, we integrate all molecular emission along each unmasked line of sight, without considering the possible presence of multiple velocity components. Finally, we apply to the moment maps the spatial mask where the dust polarization measurements are reliable as described in Sec. 3. Similarly, as for the intensity map, the line width map shown in this figure corresponds to the measured second moment, which does not consider possible spectral complexity (multiple velocity components or the presence of an outflow) or a correction for the opacity broadening. These corrections could bias our measurement, particularly in the central part of NGC 2024, as shown by the spectral decomposition results of CN and HCO in Figs. 12 and 13 in App. B.1 and B.2 respectively).
The integrated line intensity (moment 0) is shown in the left panels of Fig. 3. Both CN and HCO show the brightest emission towards the central part of NGC 2024, at the heavily dust-obscured region, as seen in Fig. 1.
@@
The second moment map (the observed line width () map) is shown in the right panels of Fig. 3. We note that to first order the HCO line is broader than CN, which has several possible reasons. Firstly, HCO is brighter and more spatially distributed than CN and, therefore, could result in having a broader line. Second, HCO emission can be optically thick (see, e.g., Barnes & Crutcher, 1990), which additionally broadens the line. Third, it is possible that HCO, similarly to CO and CO has multiple velocity layers (Gaudel et al., 2023). Nevertheless, we attempt to correct these effects in our analysis described in the following sections. The regions we analyze in this work do not contain multiple velocity features in the CN emission (see Fig. 12). However, in the case of HCO emission, we do not spectrally resolve the multiple-component HCO emission. However, we note the possibility of HCO showing multiple velocity components (Fig. 13) in NGC 2024. The more thorough analysis of the HCO velocity field is, therefore, beyond the scope of this paper, and will be an aim of the upcoming studies.
@@ 4.2 Measuring turbulent velocity dispersion
The equations presented above are crucial at sub-pc scales since these effects significantly contribute to the measured line width at these scales. For example, the optical depth effect could broaden the line by a few tens of percent, which is the case of CN and HCO. This is discussed in the following section.
@@ 4.3 Radiative transfer results
The output parameters from the RADEX modelling are the excitation temperature, and opacity, . We additionally compute the peak temperature of the spectrum generated from each model using SpectralRadex. In the case of CN, we calculate excitation temperature and the opacity of each component of the multiplet, which is scaled using information about relative intensities (Tab. 1).
@@ 5.1 Davis-Chandrasekhar-Fermi method
where is the gas volume density, is the non-thermal velocity dispersion (), where is the measured FWHM of the line, corrected for the line broadening and opacity effects - Eq. 8 and 10 - Sec. 4.2, and is the spatial dispersion of magnetic field angle.
This equation is in CGS units, and the DCF method also assumes that any perturbation in the magnetic field originates from local, small-scale turbulence. The stronger the magnetic field, the smaller will be the perturbation caused by turbulence. By calculating all constants and keeping the units of number density, line width and the angle dispersion in cm, km s, and deg, respectively (e.g., Lyo et al., 2021), Eq. 13 can be expressed as:
We have included in the above a factor of 0.5 for overestimation of the magnetic field strength due to line of sight integration effects (see, e.g., Ostriker et al., 2001).
@@ 5.2 Skalidis-Tassis method
The DCF method is widely used in the literature to derive the strength of the magnetic field (see, e.g., Pattle et al., 2019). However, it assumes that isotropic turbulence and Alfvénic waves in an incompressible medium cause the observed dispersion in polarization angle. Other mechanisms such as magneto-hydrodynamic waves (Heyvaerts & Priest, 1983) and entropy modes (Lithwick & Goldreich, 2001) also cause fluctuations in polarization angle. Therefore, to derive the magnetic field strength in NGC 2024, it is important to acknowledge the contribution of non-Alfvénic motions and the compressible nature of the ISM. In this work, we derive the using the prescription presented in Skalidis & Tassis (2021); Skalidis et al. (2021):
@@ 5.3 Sliding window
We present the dispersion of the mean direction of the magnetic field in Fig. 4. We produced this map by using a ”sliding window” to remove the impact of gradients of the large-scale magnetic field (e.g., ”unsharp-masking”, Dharmawansa et al., 2009; Pattle et al., 2017). The ”sliding window” technique is based on computing the mean and standard deviation of the magnetic field orientation within the window that is three times bigger than the beam and has pixels. The size of the sliding window is made to ensure that we remove a large-scale magnetic field contribution without losing information on the small-scale perturbation of the magnetic field (Pattle et al., 2017). As shown in Mardia & Jupp (1999), a suitable way to estimate the mean and standard deviation from the directions of each window is to compute:
@@ 5.4 Magnetic field strength in NGC 2024
The reported measurements of the angle dispersion and the POS magnetic field strength, including their uncertainties for both methods, are shown in Tab. 3. We show the comparison between the magnetic field strengths derived using the DCF and the ST method, including the use of different methods to estimate the angle dispersion in Fig. 26 in App. D.3. Overall, the POS magnetic field strength derived using the prescription from Skalidis & Tassis (2021); Skalidis et al. (2021) is lower than that derived using classical DCF. In the following, we comment on the magnetic field strengths derived using the Eq. 16 in Sec. 5.2. The strength is in the range of µG. The magnetic field strength inferred from CN is overall smaller than that derived using HCO , mainly due to HCO showing slightly larger line widths than the CN in some regions, which directly impacts the strength of the magnetic field. Due to ambipolar diffusion effects (e.g. Zweibel, 2004; Tritsis et al., 2023), it is expected that the emission lines of neutral molecular species (CN in our case) are wider than those of the ion molecular species (such as HCO) (Li & Houde, 2008; Yin et al., 2021). However, in our case, we observe that HCO is broader than the CN spectral lines, even after corrections for the opacity broadening. This result has two implications. Firstly, our observations did not spectrally resolve HCO emission, which potentially results in larger line widths caused by the blending of velocity components. Secondly, it is possible that our measurements do not probe ambipolar diffusion scale ( pc, Li & Houde, 2008) in NGC 2024.
The strongest magnetic field, derived from CN and HCO is observed in the region at the intersection of the dusty filament and the edge of the expanding HII region and at the eastern side of the bubble. This result is driven mainly by the large densities and the broader CN and HCO lines observed in these regions (see Sec. C) found in this work. The magnetic field strength measured using both CN and HCO emission is stronger towards the eastern side of the bubble than on the western edge. We found lower line widths of CN and HCO at the western side of the bubble. The weakest magnetic field is derived towards the filament, although we note the derived from HCO in the western side of the bubble and filament to be comparable given the uncertainties.
@@ 6.1 The magnetic field in NGC 2024
All these studies found that the magnetic field shows a specific structure. In particular, the line of the sight (LOS) -field strength in the central area of NGC 2024 dominated by the dusty filament (Crutcher et al., 1999) weakens from the northeast to the southwest. Matthews et al. (2002) investigated (sub)millimeter dust polarization and modelled magnetic field in NGC 2024 with two components. The first component is related to the dense and dust obscured part, where magnetic field lines follow locations of the FIR sources (right panel in Fig. 1). The second component of magnetic field is linked to gas affected by the stellar feedback. In our case, we observe regions where these two features are dominating: Band D dust polarization traces the filament in the central and southern part of NGC 2024, whereas it is impacted by the stellar feedback at eastern and the western parts at ionization front (Eastern and Western Loop in the right panel in Fig. 1).
@@
The magnetic field lines are mainly parallel to the filament in NGC 2024, which is in the agreement with results presented in other studies (e.g., Planck Collaboration et al., 2016a; Santos et al., 2016; Pillai et al., 2015; Pattle et al., 2017). However, we observe a change in the magnetic field direction at the very southern part of the filament in Fig. 2, particularly in the Band D data, and also at the northern part of NGC 2024. Such changes in the direction of the POS magnetic field can indicate a few possibilites. Firstly, changes of the magnetic field often trace star formation (Pillai, 2009; Ward-Thompson et al., 2017) due to gravitational collapse of the gas that cause the magnetic field lines to have ”pinched” structure. Moreover, the variation of the magnetic field direction can be also a consequence of changes in the column density of the gas (Alina et al., 2019).

### 07 Feb 2024 v2

@@ 2.2 IRAM 30-meter observations of Orion B
We use spectroscopic data of cyanide (CN) and formyl cation (HCO) ( transition line) to trace the UV-irradiated gas (Bron et al., 2018). CN is present in UV-illuminated edges as a photodissociation product of HCN and HNC, resulting from UV-dominated chemistry, and can be collisionally excited by electrons and neutrals (Santa-Maria et al., 2023). Therefore, the CN rotational lines remain bright in dense UV-illuminated edges, which makes them excellent tracers of the UV-dominated regions, like the one seen in NGC 2024. In addition, CN shows the multiple hyperfine components (Penzias et al., 1974), spectrally resolved by our observations. We report the properties of each hyperfine component used in this work in Tab. 1, which allow the opacity (Milam et al., 2009) and excitation temperature of the (1–0) transition to be derived, enabling its excitation to be constrained accurately and for limits on the gas density to be obtained.
HCO , similarly to CN, traces photon-dominated regions (Young Owl et al., 2000a; Lis & Schilke, 2003). HCO has different production pathways dominant at different physical conditions of the gas. At the edges of the dissociation region, HCO is a precursor of CO, and can be produced through the CO (Hogerheijde et al., 1995; Young Owl et al., 2000b; van der Werf et al., 1996) or CH molecule (Goicoechea et al., 2016). Recent studies detected HCO emission the edge of the PDR in the Orion bar (Goicoechea et al., 2016), and the Horsehead nebula (Hernández-Vera et al., 2023). Pety et al. (2017); Gratier et al. (2017) found that CN and HCO are sensitive to the UV radiation. Moreover, by applying a clustering algorithm, Bron et al. (2018) found that CN and HCO trace the UV-radiated gas, contrary to the CO molecule, which gets photodissociated. Therefore, using CN and HCO emission in our work, together with the FUV heated dust traced by SOFIA HAWC+ observations, we can directly assess the gas impacted by the radiative stellar feedback in NGC 2024.
@@ 3 Dust polarization
The dust continuum emission is strongest toward the center of NGC 2024. A significant emission is observed along the filament that is in the front of the ionization region caused by the young massive star IRS 2b (the central part of our map, Bik et al., 2003), and along the ionization fronts on the southeastern and southwestern part of NGC 2024 (labeled as Bubble West and Bubble East, and shown in the middle row in Fig. 2).
@@
In addition, we compare our magnetic field directions to the direction inferred from the near-infrared (NIR) polarization of young stars (Kandori et al., 2007). The NIR polarization directly traces the direction of the magnetic field (that is, there is no difference as for the FIR dust polarization - Sec. 2.1). We show this comparison in the top panels of Fig. 11 in App. A.2. As seen from the figure, we find that magnetic field vectors derived from NIR polarization and our work are parallel to each other, indicating an agreement between these two datasets in NGC 2024 (see Fig. 11 in App. A.2). In addition, we compare the magnetic field direction with the 100 µm (Dotson et al., 2000) and submillimeter (850 µm) polarization measurements (Matthews et al., 2002) in the central part of NGC 2024 (labeled with the dashed black rectangle in top panels of Fig. 11) and show zoomed-in panels in the bottom row of Fig. 11.
@@
The area labeled with dark blue contours in the top panels in Fig. 11 is particularly interesting because we observe an outflowing feature in the HCO emission at velocities of km s. The outflow originates from the FIR 5 source located in the dense molecular cloud behind the expanding HII region (Richer et al., 1992; Greaves et al., 2001; Choi et al., 2015). Dust continuum emission at 154 µm possibly traces the outflow in this region (see also the bottom middle panel in Fig. 3), as we note that magnetic field lines follow the direction of the outflow. We have not observed this behavior at 214 µm.
@@
The magnetic flux is frozen within the molecular gas, and as a consequence, the field will trace its morphology. Therefore, the local environmental conditions that shape the distribution of molecular gas will also impact the magnetic field. On the one hand, the magnetic field is highly ordered in some regions of NGC 2024; for instance, field follows the dusty filament to the south of NGC 2024 (see, for instance, the bottom left panel of Fig. 2), except for the very southern part of the filament (close to the bottom left corner of the top panel in Fig. 2), where the magnetic field is perpendicular to the filament. At the northern part of NGC 2024, the magnetic field direction varies from following the filament to being perpendicular.
@@
We observe the nearly horizontal direction of the -field at the edges of the expanding HII region (middle panels in Fig. 2). Gas affected by the stellar feedback (for example, UV radiation, stellar winds) is pushed outward from the HII region. Consequently, magnetic field lines become parallel to the edges of such an expanding shell (for example, Tahani et al., 2023). On the other hand, the magnetic field appears chaotic in the central and densest parts of NGC 2024. This area shows great complexity, as it results from the mixture of the filament located in the front, the HII region and the dense molecular gas in the background (see Fig. 8 in Matthews et al. 2002, or Fig. 5 in Roshi et al. 2014). One possible explanation for this observed morphology of the magnetic field is that the magnetic field changes its orientation relative to the filament depending on the column density contrast between the filament and the background emission, from being parallel to perpendicular to the molecular gas distribution ( cm, for example, Planck Collaboration et al., 2016c; Alina et al., 2019). Additionally, the magnetic field could be ”pinched” due to the gravitational collapse of the gas (Basu, 2000; Lai et al., 2002; Doi et al., 2020, 2021), which can also explain the magnetic field direction at the northern part of NGC 2024 and the very southern part of the filament.
@@ 4.1 Moment maps
Fig. 3 shows moment maps for the CN(1–0) (top row) and HCO(1–0) (bottom row) emission. We create these moment maps using CUBE in GILDAS. Before creating moment maps, we mask CN and HCO cubes. The mask is created using the segmentation technique in CUBE, where we identify neighboring voxels with continuous S/N; we have selected all voxels with S/N above 2 (see Einig et al., 2023, for a complete description). In this case, we integrate all molecular emission along each unmasked line of sight, without considering the possible presence of multiple velocity components. Finally, we apply to the moment maps the spatial mask where the dust polarization measurements are reliable as described in Sec. 3. Similarly, as for the intensity map, the line width map shown in this figure corresponds to the measured second moment, which does not consider possible spectral complexity (multiple velocity components or the presence of an outflow) or a correction for the opacity broadening. These corrections could bias our measurement, particularly in the central part of NGC 2024, as shown by the spectral decomposition results of CN and HCO in Figs. 12 and 13 in App. B.1 and B.2 respectively).
The integrated line intensity (moment 0) is shown in the left panels of Fig. 3. Both CN and HCO show the brightest emission toward the central part of NGC 2024, at the heavily dust-obscured region, as seen in Fig. 1.
@@
The second moment map (the observed line width () map) is shown in the right panels of Fig. 3. We note that to first order the HCO line is broader than CN, which has several possible reasons. Firstly, HCO is brighter and more spatially distributed than CN and, therefore, could result in having a broader line. Second, HCO emission can be optically thick (see, Barnes & Crutcher, 1990), which additionally broadens the line. Third, it is possible that HCO, similarly to CO and CO has multiple velocity layers (Gaudel et al., 2023). Nevertheless, we attempt to correct these effects in our analysis described in the following sections. The regions we analyze in this work do not contain multiple velocity features in the CN emission (see Fig. 12). However, in the case of HCO emission, we do not spectrally resolve the multiple-component HCO emission. However, we note the possibility of HCO showing multiple velocity components (Fig. 13) in NGC 2024. The more thorough analysis of the HCO velocity field is, therefore, beyond the scope of this paper, and will be an aim of the upcoming studies.
@@ 4.2 Measuring turbulent velocity dispersion
The equations presented above are crucial at subparsec scales since these effects significantly contribute to the measured line width at these scales. For example, the optical depth effect could broaden the line by a few tens of percent, which is the case of CN and HCO. This is discussed in the following section.
@@ 4.3 Radiative transfer results
The output parameters from the RADEX modeling are the excitation temperature, and opacity, . We additionally compute the peak temperature of the spectrum generated from each model using SpectralRadex. In the case of CN, we calculate excitation temperature and the opacity of each component of the multiplet, which is scaled using information about relative intensities (Tab. 1).
@@ 5.1 Davis-Chandrasekhar-Fermi method
where is the gas volume density, is the nonthermal velocity dispersion (), where is the measured FWHM of the line, corrected for the line broadening and opacity effects - Eq. 8 and 10 - Sec. 4.2, and is the spatial dispersion of magnetic field angle.
This equation is in CGS units, and the DCF method also assumes that any perturbation in the magnetic field originates from local, small-scale turbulence. The stronger the magnetic field, the smaller will be the perturbation caused by turbulence. By calculating all constants and keeping the units of number density, line width and the angle dispersion in cm, km s, and deg, respectively (Lyo et al., 2021), Eq. 13 can be expressed as:
We have included in the above a factor of 0.5 for overestimation of the magnetic field strength due to line of sight integration effects (see, for instance, Ostriker et al., 2001).
@@ 5.2 Skalidis-Tassis method
The DCF method is widely used in the literature to derive the strength of the magnetic field (see, Pattle et al., 2019). However, it assumes that isotropic turbulence and Alfvénic waves in an incompressible medium cause the observed dispersion in polarization angle. Other mechanisms such as magneto-hydrodynamic waves (Heyvaerts & Priest, 1983) and entropy modes (Lithwick & Goldreich, 2001) also cause fluctuations in polarization angle. Therefore, to derive the magnetic field strength in NGC 2024, it is important to acknowledge the contribution of non-Alfvénic motions and the compressible nature of the ISM. In this work, we derive the using the prescription presented in Skalidis & Tassis (2021); Skalidis et al. (2021):
@@ 5.3 Sliding window
We present the dispersion of the mean direction of the magnetic field in Fig. 4. We produced this map by using a ”sliding window” to remove the impact of gradients of the large-scale magnetic field (”unsharp-masking”, Dharmawansa et al., 2009; Pattle et al., 2017). The ”sliding window” technique is based on computing the mean and standard deviation of the magnetic field orientation within the window that is three times bigger than the beam and has pixels. The size of the sliding window is made to ensure that we remove a large-scale magnetic field contribution without losing information on the small-scale perturbation of the magnetic field (Pattle et al., 2017). As shown in Mardia & Jupp (1999), a suitable way to estimate the mean and standard deviation from the directions of each window is to compute:
@@ 5.4 Magnetic field strength in NGC 2024
The reported measurements of the angle dispersion and the POS magnetic field strength, including their uncertainties for both methods, are shown in Tab. 3. We show the comparison between the magnetic field strengths derived using the DCF and the ST method, including the use of different methods to estimate the angle dispersion in Fig. 26 in App. D.3. Overall, the POS magnetic field strength derived using the prescription from Skalidis & Tassis (2021); Skalidis et al. (2021) is lower than that derived using classical DCF. In the following, we comment on the magnetic field strengths derived using the Eq. 16 in Sec. 5.2. The strength is in the range of µG. The magnetic field strength inferred from CN is overall smaller than that derived using HCO , mainly due to HCO showing slightly larger line widths than the CN in some regions, which directly impacts the strength of the magnetic field. Due to ambipolar diffusion effects (Zweibel, 2004; Tritsis et al., 2023), it is expected that the emission lines of neutral molecular species (CN in our case) are wider than those of the ion molecular species (such as HCO) (Li & Houde, 2008; Yin et al., 2021). However, in our case, we observe that HCO is broader than the CN spectral lines, even after corrections for the opacity broadening. This result has two implications. Firstly, our observations did not spectrally resolve HCO emission, which potentially results in larger line widths caused by the blending of velocity components. Secondly, it is possible that our measurements do not probe ambipolar diffusion scale ( pc, Li & Houde, 2008) in NGC 2024.
The strongest magnetic field, derived from CN and HCO is observed in the region at the intersection of the dusty filament and the edge of the expanding HII region and at the eastern side of the bubble. This result is driven mainly by the large densities and the broader CN and HCO lines observed in these regions (see Sec. C) found in this work. The magnetic field strength measured using both CN and HCO emission is stronger toward the eastern side of the bubble than on the western edge. We found lower line widths of CN and HCO at the western side of the bubble. The weakest magnetic field is derived toward the filament, although we note the derived from HCO in the western side of the bubble and filament to be comparable given the uncertainties.
@@ 6.1 The magnetic field in NGC 2024
All these studies found that the magnetic field shows a specific structure. In particular, the line of the sight (LOS) -field strength in the central area of NGC 2024 dominated by the dusty filament (Crutcher et al., 1999) weakens from the northeast to the southwest. Matthews et al. (2002) investigated (sub)millimeter dust polarization and modeled magnetic field in NGC 2024 with two components. The first component is related to the dense and dust obscured part, where magnetic field lines follow locations of the FIR sources (right panel in Fig. 1). The second component of magnetic field is linked to gas affected by the stellar feedback. In our case, we observe regions where these two features are dominating: Band D dust polarization traces the filament in the central and southern part of NGC 2024, whereas it is impacted by the stellar feedback at eastern and the western parts at ionization front (Eastern and Western Loop in the right panel in Fig. 1).
@@
The magnetic field lines are mainly parallel to the filament in NGC 2024, which is in the agreement with results presented in other studies (such as, Planck Collaboration et al., 2016a; Santos et al., 2016; Pillai et al., 2015; Pattle et al., 2017). However, we observe a change in the magnetic field direction at the very southern part of the filament in Fig. 2, particularly in the Band D data, and also at the northern part of NGC 2024. Such changes in the direction of the POS magnetic field can indicate a few possibilites. Firstly, changes of the magnetic field often trace star formation (Pillai, 2009; Ward-Thompson et al., 2017) due to gravitational collapse of the gas that cause the magnetic field lines to have ”pinched” structure. Moreover, the variation of the magnetic field direction can be also a consequence of changes in the column density of the gas (Alina et al., 2019).